OpenStudy (anonymous):

show that the equation 6x^4-7x+1=0 does not have more than two distinct reals roots (use Roll's Theorem.)

OpenStudy (amistre64):

isnt rolls thrm just the mean value thrm at zero?

OpenStudy (amistre64):

how many solutions to dy/dx = 0 ?

OpenStudy (amistre64):

6x^4-7x+1=0 24x^3 -7 = 0 x^3 = 7/24 since there is only one value of dy/dx that equals zero, there must be at most 2 real roots